Qs=-x^2 +600x +25


1
Expert's answer
2021-11-10T18:12:39-0500

"-x^2 +600x +25=0"


"x=\frac{600\pm \sqrt{360000+100}}{2}"


"x_1=205,x_2=395"


"Q_s=-(x-205)(x-395)"